# DAG-Plan: Generating Directed Acyclic Dependency Graphs for Dual-Arm Cooperative Planning

### v2

@@ 4.1 Experimental Setup
To validate the correctness and execution efficiency of our method, we created a Dual-arm Kitchen Benchmark, including plan tests and physical simulation tests. This benchmark fills the gap in long-sequence operation rigid body physics simulation benchmarks for dual-arm robots. The goal of this benchmark is to validate the success rate and efficiency of dual-arm robot planning in complex scenarios. The dual-arm robot should correctly plan and fully utilize both the left and right arms, completing tasks with as few execution stages as possible. The benchmark consists of 9 sequential tasks (see details in Appendix B) are shown in Table 1, comprising a total of 78 sub-tasks. The main setting is a modern kitchen divided into three scenarios shown in Figure 3, containing a total of 26 assets and objects. Our physical simulation scene is built on the Sapien [21]. The embodied platform for planning execution is a dual-arm robot with a mobile base, each 7-degree-of-freedom arm equipped with a two-finger gripper. The platform is equipped with five RGB-D cameras, attached above and below each arm’s gripper and on the robot’s head.
Table 1: Task list of Dual-arm Kitchen Benchmark.
Index Instruction Index Instruction
Task 1 Put the apple and bread onto the plate Task 6 Heat the soup and pour a cup of cola
Task 2 Place the apple on the plate and toast the bread Task 7 Put the apple and the pear into the bowl
Task 3 Juice the apple and toast the bread Task 8 Cut the carrot and red bell pepper
Task 4 Wash the cup and the bowl Task 9 Make a pot of soup
Task 5 Heat the soup and put the tin on the table
Figure 3: Snapshots of 3 Scenarios of Dual-arm Kitchen Benchmark. Scenario 1 encompasses tasks 1-3, scenario 2 encompasses tasks 4-6, and scenario 3 encompasses tasks 7-9.
@@
Evaluation of Planning Effectiveness and Conciseness. In this experiment, we focus on testing the conciseness of the generated plans and the number of stages required. In a stage, the robot can execute a right arm node and a left arm node. This requires that the plans generated by the LLMs can achieve the task goals in terms of language logic and do not violate the preconditions for stage execution. Since the lower-level execution method uses reinforcement learning, learning from incorrect plans is costly and unproductive, making plan validation crucial. We used GPT-4111https://openai.com/api. This work mainly uses gpt-4-turbo-2024-04-09. under the temperature of 0.5 to generate five plans for each task, evaluating their Success Rate (SR) and the fewest number of Stage of the passed plan required at the language level. Finally, we calculated the average success rate and the average number of stages for all tasks. We defined Stage Efficiency as the ratio of single-arm plan stages to the stages required by each method. For failed dual-arm plans, we calculated the stage count based on the single-arm plan stages to ensure a fair comparison.
@@
Evaluation with Physical Simulation. In this experiment, we will test the executability and execution efficiency of the plans in physical simulation scenarios. Compared to plan tests, physical simulation tests validate both the high-level planning and low-level execution capabilities. We will use reinforcement learning to acquire executable skills from the plans that passed the plan tests and then execute them in the physical simulation environment. We select the plan that passes the plan test and requires the minimum stages for skill learning. Then evaluate the Success Rate (SR) of execution and minimum execution Time in the physical environment with 10 trials. Finally, we calculated the average success rate and the execution time for all tasks. We defined Execution Efficiency as the ratio of single-arm plan execution time to the execution time required by each method. For failed dual-arm plans, we calculated the average time based on the single-arm execution time to ensure a fair comparison.
@@
Task Planning for Single-arm (TP-S) directly uses LLMs to generate a full task list, with each stage involving a single arm or both arms to manipulate a single object. Task Planning for Dual-arm (TP-D) also directly uses LLMs to generate a full task list, but each stage can use the arms to manipulate either a single object or two different objects. Our method, DAG-Plan, generates a task graph, followed by task planning inference to iteratively generate nodes for each stage.
4.2 Experimental Results
Evaluation of Planning Effectiveness and Conciseness. As shown in Table 2, in the plan tests, DAG-Plan consistently outperformed both TP-S and TP-D, showcasing superior efficiency and robustness. DAG-Plan achieved a high success rate across all tasks, demonstrating its effectiveness in dual-arm manipulation. Notably, it maintained an impressive macro average success rate of 97.8% and exhibited a significant reduction in the required stages for task completion, averaging 5.67 stages. Furthermore, the completeness of the DAG generation is high, with only an incomplete DAG generated in task 8. By re-generating it through reflection, a complete DAG was obtained, increasing the success rate from 0.8 to 1.0. These results underscore the effectiveness of DAG-Plan in achieving task goals and its efficiency in execution. In contrast, TP-S, primarily focused on the single-arm plan, generally required more stages to complete tasks compared to TP-D and DAG-Plan. Although TP-S maintained a relatively high and consistent success rate, it was less efficient in stage minimization. Moreover, TP-D, relying on language models to generate dual-arm task plans, exhibited a significantly lower success rate (macro average SR of 44.5%), and often produced plans that were not executable in the physical environment. While theoretically capable of reducing the number of stages required for tasks, TP-D frequently encountered challenges related to coordination complexities and unrealistic task assignments. This highlights the superiority of DAG-Plan in effectively translating high-level plans into physical actions and navigating complexities in the environment with relative ease compared to TP-D. Additionally, we provide detailed plans of DAG-Plan and the baseline in Appendix C.
Table 2: Performance comparison on plan tests. We report the success rate and minimum stage of the 5 plans generated by the LLM for each task, and the macro average metrics for all tasks.
Scenario1 Scenario2
Task1 Task2 Task3 Task4 Task5 Task6
SR Stage SR Stage SR Stage SR Stage SR Stage SR Stage
TP-S $1.0$ $4$ $0.8$ $6$ $1.0$ $8$ $1.0$ $8$ $1.0$ $10$ $0.8$ $11$
TP-D $1.0$ $2$ $0.0$ $Fail$ $1.0$ $4$ $0.0$ $Fail$ $0.2$ $6$ $0.6$ $6$
DAG-Plan $1.0$ $2$ $1.0$ $4$ $1.0$ $4$ $0.8$ $6$ $1.0$ $6$ $1.0$ $6$
Scenario3 Macro Avg
Task7 Task8 Task9 SR Stage Stage Efficiency
SR Stage SR Stage SR Stage
TP-S $1.0$ $8$ $0.6$ $10$ $1.0$ $13$ $91.1\%$ $8.67\pm 2.54$ $100.0\%$
TP-D $0.6$ $5$ $0.6$ $7$ $0.0$ $Fail$ $44.5\%$ $6.33\pm 2.87$ $137.0\%$
DAG-Plan $1.0$ $5$ $1.0$ $7$ $1.0$ $11$ $97.8\%$ $5.67\pm 2.35$ $152.9\%$
Table 3: Performance comparison on physical simulation tests. We report the success rate and minimum time for each task with 10 trials, and the macro average metrics for all tasks.
Scenario1 Scenario2
Task1 Task2 Task3 Task4 Task5 Task6
SR Time SR Time SR Time SR Time SR Time SR Time
TP-S $0.7$ $37.1$ $0.7$ $59.0$ $0.7$ $79.6$ $1.0$ $84.1$ $0.3$ $105.5$ $0.3$ $114.4$
TP-D $1.0$ $18.6$ $0.0$ $Fail$ $0.6$ $40.1$ $0.0$ $Fail$ $0.0$ $Fail$ $0.0$ $Fail$
DAG-Plan $1.0$ $18.6$ $0.7$ $39.4$ $0.6$ $40.1$ $1.0$ $66.3$ $0.3$ $63.2$ $0.3$ $76.3$
Scenario3 Macro Avg
Task7 Task8 Task9 SR Time Execution Efficiency
SR Time SR Time SR Time
TP-S $0.5$ $74.0$ $0.1$ $105.2$ $0.2$ $136.2$ $49.9\%$ $88.3\pm 28.5$ $100.0\%$
TP-D $0.5$ $55.1$ $0.3$ $76.4$ $0.0$ $Fail$ $26.7\%$ $76.6\pm 35.5$ $115.3\%$
DAG-Plan $0.6$ $53.3$ $0.3$ $76.4$ $0.2$ $107.4$ $55.6\%$ $60.1\pm 24.5$ $147.0\%$
Evaluation with Physical Simulation. As shown in Table 3, the physical simulation tests provided further insights into the practical applicability and execution capabilities of our planning methods under more dynamic and realistic conditions. Here again, DAG-Plan demonstrated a balanced performance with a solid success rate and efficient execution times. We show the execution process of DAG-Plan in a physical simulation environment in Figure 4. Moreover, we provide detailed analysis and explanation of DAG-Plan and the baseline in Appendix D.
Compared to TP-D, DAG-Plan effectively translates high-level plans into feasible actions based on target object information and the robot’s current state under the guidance of a task graph. Both DAG-Plan and TP-S were able to complete $/9$ 9 tasks in the physical simulation tests. However, while TP-D passed the plan tests for $6/9$ tasks, it only completed $4/9$ in the physical simulation tests. In tasks 5 and 6, the plans generated by TP-D could not be executed due to real-world physical constraints. For example, in task 6, the plan of TP-D included “put cola bottle into refrigerator cooler" with left arm and “close microwave door" with right arm. However, since the refrigerator is on the right side of the microwave, the dual-arm robot could not cross its arms to execute this action. Ultimately, DAG-Plan achieved a 28.9% higher success rate in physical simulation tests compared to TP-D. This demonstrates that, both in high-level planning and low-level execution, the success rate of the dual-arm plans generated by TP-D is significantly inferior to those generated by DAG-Plan.
Regarding execution efficiency, the dual-arm plan allows for parallel execution of sub-tasks, resulting in higher efficiency. DAG-Plan’s execution efficiency was 47.0% higher than TP-S, as it maximized the parallelization of sub-tasks while ensuring the feasibility of the plan. In all tasks, DAG-Plan’s execution time was shorter than TP-S. Although TP-D had similar execution times to DAG-Plan for some tasks, its overall efficiency was low because $5/9$ tasks could not be executed. Consequently, TP-D’s efficiency was only 15.3% higher than TP-S.
Overall, the experimental results strongly support the adoption of DAG-Plan in complex robotic operations. DAG-Plan not only outperforms traditional single-arm and dual-arm planning approaches in terms of success rates and efficiency but also demonstrates significant robustness and reliability in translating plans into actionable steps in a physical context.
@@ 4.2 Experimental Results
Figure 4: Head camera snapshots of the execution process of 2 example long-horizon tasks.
5 Conclusion
This work introduces DAG-Plan, which efficiently and accurately generates collaborative plans with LLMs for mobile dual-arm robots. DAG-Plan decomposes complex tasks into directed acyclic graph (DAG) with clear temporal relationships and iteratively selects feasible sub-tasks based on environmental observations during execution. The main contribution is the replacement of task sequences with a DAG and the dynamic adjustment of the planning according to the current situation, allowing dual-arm robots to flexibly utilize both arms for sub-task execution. We also conducted a dual-arm kitchen benchmark, providing a testing scenario for future long-horizon dual-arm works.
Limitations and Future Works. Currently, DAG-Plan relies on reinforcement learning for underlying skills, which may be inefficient for complex tasks. Our future efforts will aim to enhance skill learning modules, boosting automation and success rates for real-world applications.
Appendix A Implementation Details of Motion Planning and Reinforcement Learning
A.1 Embodied Platform
We used the Baxter robot from Rethink Robotics as the primary embodied platform shown in Figure 5. We configured its base in the simulation as a mobile base capable of translating on the plane but not rotating. The grippers of the Baxter robot were replaced with Robotiq 85 grippers to achieve more robust grasping.
Figure 5: Embodied platform of DAG-Plan.
A.2 Motion Planning
For motion planning, we use open-source implementations of RRT-connect [^2] https://github.com/haosulab/mplib. We manually set multiple candidate poses and use motion planning to find the pose that requires the shortest movement steps while avoiding collisions. For sub-tasks of the types occupy, tool use, and operate, we use motion planning to move the end-effector to within 0.15 cm of the target object, facing it as directly as possible. For release type sub-tasks, we adjust based on the opening direction of the target object. For high-precision sub-tasks that require both arms, such as “put pot into sink”, where motion planning alone struggles to maintain fixed distance and equal height for stable placement, we directly use reinforcement learning to learn the sub-task.
A.3 Reinforcement Learning
For reinforcement learning training, we use open-source implementations of PPO [^3] https://github.com/DLR-RM/stable-baselines3 algorithm, and list the hyper-parameters in Table 4. The observation space is the low-level state of the objects and robot and the images captured by 5 RGB-D cameras. The control mode of mobile base adopts velocity control. The control mode of the dual-arm adopts arm joint delta position control, which means the action space consists of the change of the arm joint position.
Table 4: Hyper-parameter of PPO algorithm.
Hyper-parameter Value
Discount factor $\gamma$ $0.99$
# of epochs per update $10$
@@ A.3 Reinforcement Learning
# of environments $32$
Batch size $400$
Target KL divergence None
@@
# of steps per update $4800$
Rollout steps per episode $100$
Corresponding rollout seconds per episode $4s$
Shape of observation space $(128,128,20)$Visual input
$95$State input
Shape of action space $18$
In this work, we use a set of reward primitives shown in Table 5. We use LLMs to combine reward primitives and specify the corresponding object and target object, with manual adjustments for corrections to generate dense rewards. The reward function and success condition for each sub-task are shown in Table 6. In the final reward, we normalize the reward functions of the subtasks corresponding to the left and right hands to 1 and assign additional rewards for successful states. The reward $r$ is represented as:
$r=r_{\text{right}}+r_{\text{left}}+r_{\text{side\_success}}\left\{\begin{aligned} &30,\text{if side success first time}\\ &0,\enspace\text{otherwise}\end{aligned}\right.+r_{\text{success}}\left\{\begin{aligned} &150-\text{steps},\text{if success}\\ &0,\quad\quad\quad\quad\enspace\text{otherwise}\end{aligned}\right.$
Table 5: Reward primitives.
Reward Primitives Description Pseudocode
$obj\_dis$ Move object close to target object $\frac{1}{2}(1-\frac{\text{dis}(obj,target)}{init\_dis}+1-\text{tanh}(\text{dis}(obj,target)))$

### v3

@@ IV-A Experimental Setup
To validate the correctness and execution efficiency of our method, we created a Dual-arm Kitchen Benchmark, including plan tests and physical simulation tests. This benchmark fills the gap in long-sequence operation physics simulation benchmarks for dual-arm robots. The goal of this benchmark is to validate the success rate and efficiency of dual-arm robot planning and executable ability in complex scenarios. The dual-arm robot should correctly plan and fully utilize both the left and right arms, completing tasks with as few execution stages as possible. The benchmark consists of 5 sequential tasks are shown in Table I and Figure 3, comprising a total of 44 sub-tasks. Our physical simulation scene is built on the Sapien [34]. The embodied platform is Agilex CobotMagic, each 6 DOF arm equipped with a two-finger gripper. The platform is equipped with an Intel RealSense L515 RGB-D camera, attached on the robot’s head.
IV-A1 Evaluation of Planning Effectiveness and Conciseness
@@
In this experiment, we focus on testing the conciseness of the generated plans and the number of stages required. In a stage, the robot can execute a right arm node and a left arm node. This requires that the plans generated by the LLMs can achieve the task goals in terms of language logic and do not violate the preconditions for stage execution. We used LLMs to generate 10 plans for each task, evaluating their Success Rate (SR) and the minimum Stage of the passed plan required at language level. Due to cost considerations, we also include average Tokens usage as an important metric for evaluating cost of the algorithm. Finally, we calculated the average success rate and the average number of stages for all tasks. We defined Stage Efficiency as the ratio of single-arm plan stages to the stages required by each method. For failed dual-arm plans, we calculated the stage count based on the single-arm plan stages to ensure a fair comparison.
IV-A2 Evaluation with Physical Simulation
@@
In this experiment, we will test the executability and execution efficiency of the plans in physical simulation scenarios. Compared to plan tests, physical simulation tests validate both the high-level planning and low-level execution capabilities. Then evaluate the Success Rate (SR) and minimum Time which consists of query time and execution time. in the physical environment with 10 trials. Finally, we calculated the average success rate and the time for all tasks. We defined Execution Efficiency as the ratio of single-arm plan time to the time required by each method. For failed dual-arm plans, we calculated the average time based on the single-arm time to ensure a fair comparison.
IV-A3 Baseline Algorithms and Proposed Method
We compare the planning algorithms baselines and our proposed method:
  1. 1.
@@ IV-A3 Baseline Algorithms and Proposed Method
    Task Planning for Single-arm (TP-S): TP-S directly uses LLMs to generate a full task sequence, with each stage involving a single arm to manipulate a single object.
  2. 2.
    Task Planning for Dual-arm (TP-D): TP-D directly uses LLMs to generate a full task sequence, with each stage involving dual arms.
  3. 3.
    Twostep [11]: Twostep decomposes complex tasks into fully independent sub-tasks with dual agents, with the primary agent and the secondary agent each completing separate sub-tasks to accelerate task execution.
  4. 4.
    RoCo [12]: RoCo proposed multi-agents collaboration using LLMs for planning which is an iterative method. Robots query LLMs in each stages to collaboratively reason about task strategies. DABICO [13] is a dual-armed version of the RoCo, and is equivalent to the RoCo in our experiments settings.
  5. 5.
    DAG-Plan (Proposed): Our method DAG-Plan generates a task graph, followed by task planning inference to iteratively generate nodes for each stage.
IV-B Evaluation of Planning Effectiveness and Conciseness
@@
Figure 4: Simulation snapshots of the execution process of long-horizon task 2.
TABLE II: Performance comparison on plan tests. We report the success rate, minimum stage and average token usage of the 10 plans generated by the LLM for each task.
Task1 Task2 Task3 Task4 Task5 Average
@@
TP-S $100\%$ $8$ $2030.0$ $70\%$ $11$ $2424.8$ $100\%$ $6$ $2086.8$ $90\%$ $12$ $2413.0$ $100\%$ $7$ $2096.6$ $92\%$ $100.0\%$ $2210.2$
@@
TwoStep[11] $100\%$ $4$ $3699.6$ $0\%$ $Fail$ $4248.2$ $0\%$ $Fail$ $3823.4$ $0\%$ $Fail$ $4130.2$ $100\%$ $5$ $3810.2$ $40\%$ $115.8\%$ $3942.3$
RoCo[12] $100\%$ $4$ $26615.0$ $0\%$ $Fail$ $68921.4$ $100\%$ $4$ $25410.2$ $40\%$ $8$ $65242.0$ $100\%$ $4$ $26985.0$ $68\%$ $141.9\%$ $42634.7$
DAG-Plan $100\%$ $4$ $2031.0$ $70\%$ $7$ $2207.2$ $100\%$ $4$ $1980.6$ $70\%$ $7$ $2277.0$ $100\%$ $4$ $1973.0$ $88\%$ $169.2\%$ $2093.8$
TABLE III: Performance comparison on physical simulation tests. We report the success rate and minimum time for each task with 10 trials. Time consists of query time and execution time.
Task1 Task2 Task3 Task4 Task5 Average